OpenStudy (anonymous):

The probability of rain on Thanksgiving Day in Boston is 0.65. The probability of rain in Honolulu on Thanksgiving Day is 0.30. What is the probability that it rains in Boston or Honolulu on Thanksgiving Day? Make sure you only account for the scenario in which both Boston and Honolulu get rain on Thanksgiving Day once.

OpenStudy (anonymous):

\[.65+.30-.65\times .30\]

OpenStudy (anonymous):

assuming these events are independent, which is probably why they had them so far apart, then \[P(A\cup B)=P(A)+P(B)-P(A)\times P(B)\]

OpenStudy (anonymous):

these questions really all over the map probability summation algebra what the hell kind of class is this? fruit salad math?

OpenStudy (anonymous):

yes! I didnt think any one would know of this class, but here we are

OpenStudy (anonymous):

0.09 @satellite73

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

compute this \[.65+.30-.65\times .30\]

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=.65%2B.3-.65*.3

OpenStudy (anonymous):

math fruit salad in the late summer no wonder it is @summertimesadness

OpenStudy (anonymous):

haha yupp